### v1

# LLM & HPC:Benchmarking DeepSeek’s Performance in High-Performance Computing Tasks

Abstract
Large Language Models (LLMs), such as GPT-4 and DeepSeek, have been applied to a wide range of domains in software engineering. However, their potential in the context of High-Performance Computing (HPC) much remains to be explored. This paper evaluates how well DeepSeek, a recent LLM, performs in generating a set of HPC benchmark codes: a conjugate gradient solver, the parallel heat equation, parallel matrix multiplication, DGEMM, and the STREAM triad operation. We analyze DeepSeek’s code generation capabilities for traditional HPC languages like C++, Fortran, Julia and Python. The evaluation includes testing for code correctness, performance, and scaling across different configurations and matrix sizes. We also provide a detailed comparison between DeepSeek and another widely used tool: GPT-4. Our results demonstrate that while DeepSeek generates functional code for HPC tasks, it lags behind GPT-4, in terms of scalability and execution efficiency of the generated code.
Keywords LLM  $\cdot$ DeepSeek  $\cdot$ GPT-4  $\cdot$ HPC  $\cdot$ Code Performance  $\cdot$ Programming language processing
1 Introduction
Large Language Models (LLMs) have rapidly advanced, potentially approaching Artificial General Intelligence (AGI). Trained on vast textual data, LLMs have shown remarkable capabilities in natural language processing and visualization tasks [1, 12, 22]. This trend has accelerated with the release of the DeepSeek models by the Chinese company DeepSeek [5]. Their DeepSeek-V3 [17, 14] was trained in two months at a cost of \$5.6 million, significantly cheaper than similar models. On January 20th, they launched DeepSeek-R1 [13], a reasoning model exhibiting performance comparable to OpenAI’s O1 model, available for open research [9]. DeepSeek-R1 is trained through large-scale reinforcement learning (RL) without supervised fine-tuning, demonstrating powerful reasoning behaviors [13].
In the high-performance computing (HPC) domain, LLMs are being explored for tasks such as code analysis, generation, and optimization. However, there is a lack of standardized, reproducible evaluation processes for LLMs in HPC-specific tasks. For this reason, this paper explores the performance of the open-source LLM DeepSeek for simple code generation tasks in HPC. We evaluate code generation using four programming languages (PLs) selected for their prominence in HPC and based on the TIOBE Index [24]. The PLs are C++ for performance and simulation, Fortran for numerical and array optimizations, Python for flexibility and integration with other HPC languages, and Julia for scientific computing performance with a Python-like syntax.
We present five examples to assess the model performance. The first example in this study features a conjugate gradient solver that incorporates matrix and vector operations. The second example is a parallel 1D stencil-based heat equation solver [7], chosen for its relevance in testing parallel computational workflows and its widespread use in scientific computing. The third example is parallel matrix multiplication, a fundamental operation in scientific computing and machine learning, often accelerated using HPC systems to enhance computational speed. The fourth example is parallel double-precision general matrix multiply (DGEMM), a widely used operation in numerical linear algebra that is essential for many scientific and engineering applications, testing the system’s computational efficiency for large-scale matrix operations. The final example is the STREAM triad operation, a key benchmark for testing memory bandwidth and data movement between CPU and memory, commonly used to evaluate the memory subsystem’s performance in HPC systems. We summarize the main contributions of this paper as follows:
  • A comparison of DeepSeek and GPT-4 for HPC code generation, specifically targeting five simple tasks.
  • A systematic evaluation of the performance and quality of code generated by DeepSeek in when writing code for four widely used HPC programming languages: C++, Fortran, Python, and Julia.
  • A comprehensive analysis of the models’ ability to generate optimized code for common HPC operations, offering insights into the potential of LLMs for high-performance computing tasks.
Table 1: Summary of LLM-based HPC related work
Paper Main Focus LLM Model Used
LLM4HPC [4] HPC Adaptation LLaMa-2
LLM4VV [20] OpenACC Testing GPT-4, CodeLlama
HPC-GPT [8] AI Model Management LLaMa-13B
Tokompiler LLM [15] Code Completion GPT-3
HPC-Coder [21] OpenMP & MPI Handling DeepSpeed
Dataset for OpenMP Translation [16] Code Translation LLaMa-2
LLMs in HPC [3] LLM-HPC Challenges Various LLMs
Godoy et al. (2023) [10] Kernel Code Generation LLaMa-2
Valero-Lara et al. (2023) [25] LLaMa-2 Comparison LLaMa-2
Godoy et al. (2024) [11] Code Parallelization GPT-3
chatHPC [26] HPC Chatbot GPT-like (based on StarCoder)
2 Related Work
LLMs have been widely used across various domains, including natural language processing (NLP) and visualization. However, their application in analyzing and optimizing HPC tasks remains challenging. In Table 1, we present the related works that use LLMs for HPC tasks. LLM4HPC [4] represents the first effort to adapt LLMs specifically for the HPC domain. The LLM4HPC framework is specifically designed for HPC, showing success in code similarity analysis, parallelism detection, and OpenMP question-answering tasks. LLM4VV [20] is a fine-tuned model that leverages the capabilities of GPT-4 and CodeLlama (based on Llama-2) to successfully generate OpenACC directives. HPC-GPT introduced by Ding et al. [8], was built on LLaMa-13B. HPC-GPT has been successfully applied to managing AI models and datasets, as well as detecting data races. Kadosh et al. [15] introduced the domain-specific Tokompiler LLM, which outperforms a GPT-3-based model in code completion and semantics for Fortran, C, and C++ code. Nichols et al. [21] present HPC-Coder. Their work demonstrated varying success in code completion, including handling OpenMP pragmas and MPI calls. Lei et al. [16] introduced a dataset designed for fine-tuning models on OpenMP Fortran and C++ code translation. Their fine-tuned model yielded more accurate results compared to GPT-4. Chen et al. [3] provide an insightful overview of the challenges and opportunities at the intersection of LLMs and HPC, extending beyond code generation. Notably, the works by Godoy et al. [10] and Valero-Lara et al. [25], which evaluate HPC kernel code generation and results for LLaMa-2, are among the first to apply LLM-based code generation to the domain of HPC software development. Additionally, Godoy et al. [11] apply LLM capabilities of GPT-3 targeting HPC kernels for code generation, and auto-parallelization of serial code in C++, Fortran, Python and Julia. Yin et al. [26] proposed chatHPC, a chatbot for HPC question answering and script generation.
3 Methodology
We used five examples: (1) a conjugate gradient solver, (2) a parallel one-dimensional heat equation solver using finite differencing, (3) parallel matrix multiplication, (4) parallel DGEMM, and (5) the STREAM triad operation. The complexity of the code increases with each example. We generated the code for these examples using DeepSeek-R1 and compared it with our previous results [6] obtained using GPT 4.0 for the first two examples. The code was generated on 02/03/2025. The queries used for code generation are shown in Table 2.
Table 2: Queries for HPC tasks code generation, whether the language was C++, Fortran, Julia, Python
Ex. Problem Query
1 Conjugate Gradient Solver Write a language code to solve the linear equation system using the conjugate gradient solver and validate it.
2 Parallel 1D Stencil-Based Heat Equation Solver Write a parallel language code to solve the one-dimensional heat equation using a finite difference scheme for the discretization in space and the Euler method for time integration, validate it and plot the solution.
3 Parallel Matrix Multiplication Write a parallel language code for matrix multiplication and validate it.
4 Double-Precision General Matrix Multiplication Write a parallel language code to perform DGEMM on large matrices, optimize the implementation for performance using parallel computing techniques, validate the results, and compare the performance with different matrix sizes and parallelization strategies.
5 STREAM Write a parallel language code to perform the STREAM triad operation on large arrays.
Example 1 : An advanced example from numerical methods textbooks is using a conjugate gradient solver to solve a system of linear equations [23].
$A^{n\times n}\cdot x^{n}=b^{n}\quad\text{with}\quad n\in\mathbb{Z}^{+},A=A^{t},\text{ and }x^{t}Ax>0,\forall X\in\mathbb{R}^{n}\text{.}$ (1)
To evaluate the code, we asked DeepSeek to use the following equation system $M\cdot x=b$ with
$M=\left(\begin{matrix}4&-1&0\\ -1&4&-1\\ 0&-1&4\end{matrix}\right),b=\left(\begin{matrix}1.0\\ 2.0\\ 3.0\end{matrix}\right)$ (2)
And the correct result is $x_{\text{exact}}=\left(\begin{matrix}13.0/28.0\\ 6.0/7.0\\ 27.0/28.0\end{matrix}\right)$.
Example 2 : Here, we want to evaluate whether DeepSeek can write the code that can solve:
$\frac{\partial u}{\partial t}=\alpha\frac{\partial^{2}u}{\partial x^{2}},\quad 0\leq x<L,t>0$ (3)
where $\alpha$ is the material’s diffusivity. For the discretization in space, a finite difference scheme was used
$u(t+1,x_{i})=u(t,x_{i})+dt\ \alpha\frac{u(t,x_{i-1})-2u(t,x_{i})+u(t,x_{i+1})}{h^{2}}$ (4)
We did not specify how to generate the grid, i.e. equidistant nodal spacing with $n$ grid points $x=\{x_{i}=i\cdot h\in\mathbb{R}|i=0,\ldots,n-1\}$, nor what time integration method to use, e.g. the Euler method.
Example 3: In this example, we evaluate DeepSeek’s ability to generate parallel code for matrix multiplication, a fundamental operation in numerical simulations and machine learning. Matrix multiplication is a critical task in scientific computing, particularly in HPC, where large matrices are multiplied to generate results in parallel. The task involves writing efficient parallel code to split matrix multiplication across multiple threads or processors, optimizing for memory usage and computational speed. The generated code is supposed to compute the following:
$C(i,j)=C(i,j)+A(i,k)*B(k,j)\text{,}\quad A,B,C\in\mathbb{R}^{n\times n}\text{.}$ (5)
Example 4: The fourth example is more general and focuses on DGEMM, a well known operation in scientific computing and HPC. The objective here is to analyze the ability of DeepSeek to generate parallel code for performing DGEMM on large-scale matrices, with a particular focus on optimizing for performance using parallel computing techniques. The generated code is supposed to compute the following:
$C=\alpha\cdot A\cdot B+\beta\cdot C,\quad A,B,C\in\mathbb{R}^{n\times n},\alpha,\beta\in\mathbf{R}\text{.}$ (6)
Example 5: The final example evaluates DeepSeek’s ability to implement the STREAM triad operation, which is commonly used in HPC benchmarks to measure memory bandwidth and data throughput. The task requires writing parallel code that performs the STREAM triad operation:
$A[i]=B[i]+scalar\cdot C[i],\quad A,B,C\in\mathbb{R}^{n},scalar\in\mathbb{R}\text{.}$ (7)
The performance of the implementation is evaluated based on its ability to efficiently utilize memory bandwidth, especially in multi-core or distributed HPC systems. The goal is to assess how well the code is generated, how well it scales as the array sizes grow, and how effectively it can handle large data sets while maintaining high computational throughput.
We copied the generated code to the paper’s GitHub repository [^1] https://github.com/NoujoudNader/AiCode_DeepSeek. For some of the generated codes, DeepSeek provided instructions on how to compile the code and some examples of expected output. These instructions were added to the Github repository.
4 Quality of the generated code
All of the generated codes were checked for compilation errors, runtime errors, and correctness. Table 3 summarizes the evaluation of all examples. For the conjugate gradient and parallel heat equation solver, the generated code with the Deep Seek model is compared with the generated code with ChatGPT 4.0 from the author’s previous work [6].
Table 3: Results for the generated code performance across two examples (Conjugate Gradient and Parallel Heat Equation). We verify that the code compiled successfully for C++and FORTRAN. We verified that the codes executed without any runtime errors, and that the code produced correct results for all languages. A comparison was made between DeepSeek and our previous results with GPT [6].
Model Deep Seek ChatGPT [6]
Language C++ Fortran Python Julia C++ Fortran Python Julia
Conjugate gradient
Compile x
Execution x
Correctness
Parallel heat equation
Compile x x
Execution x x x
Correctness x
4.1 Conjugate gradient
For Python, the generated code executed and produced the correct results. The C++ code compiled, executed, and produced the correct result. The Fortran code did not compile due to the following Error: Unexpected data declaration statement at (1) since the exact result was declared in the third last line. Moving the declaration to the top fixed the issue. After that the code compiled, executed, and produced the correct results. In Python, the exact solution was computed using np.linalg.solve from the NumPy package, and in the Julia code used x_exact = A \ b. In the other codes the exact solution was hard-coded. The Deep Seek model had one compilation issue in the Fortran code and GPT had no compilation issues. The Julia code had runtime issues for GPT but not for Deep Seek. To summarize, for the conjugate gradient method, each model had one issue with the generated code.
4.2 Parallel heat equation solver
The C++ and Python code worked without issue. The Fortran code did not compile since the variable pi was used but not declared. After declaring the variable the code worked. The Julia code had the following error ERROR: LoadError: UndefVarError: ‘@printf‘ not defined in ‘Main‘. After adding using Printf the code worked. The C++ and Fortran code used OpenMP for parallelism. The Python code used numba and Julia used Base.Threads for parallelism. The Fortran code did not compile for DeepSeek but did compile for ChatGPT. The Fortran code produced the correct result for Deep Seek but not for ChatGPT. For both models, the Julia code had runtime errors. The Python code had runtime errors for ChatGPT but no errors for DeepSeek. The model added the following parameters: heat coefficient $\alpha=0.1$, length of the bar $L=1$, nodal spacing $h=0.1$, final time $T=1$, and time step width $dt=10^{-3}$.
4.3 Parallel matrix multiplication
The matrices were filled with random numbers and the serial and parallel computation were compared. The C++, Fortran, and Python codes worked and had correct results with respect to the serial execution. The Julia code had one error while printing the results. The C++ and Fortran code used OpenMP, the Python code used the multiprocessing package, and the Julia code used Base.Threads package.
4.4 DGEMM
The Fortran code did not compile due to passed REAL(4) to REAL(8) and the result for the parallel implementation was incorrect. The Python code stopped execution with numba.core.errors.UnsupportedRewriteError. After investigating the error, we discovered that the model generated the code using nb.prange(0, n, block_size), however, the Python package numba does not provide these arguments. To fix the code, we needed to edit the numba API calls. The C++ and Fortran code used OpenMP and Julia used the package Base.Threads.
4.5 STREAM
The C++ code did not compile due to SIMD errors error: ‘c’ in ‘aligned’ clause is neither a pointer nor an array nor a reference to pointer or array. The Fortran code did not compile due to Error: Unclassifiable statement at (1). The Python code stopped with the error AttributeError: ’c_double_Array_100000000’ object has no attribute ’get_obj’. The Julia code had the following error ERROR: LoadError: UndefVarError: ‘nthreads‘ not defined in local scope. The C++ code used OpenMP, the Fortran code used coarray, the Python code used the multiprocessing package, and Julia used the Base.Threads package.
Table 4: Results for the performance of the generated code across three HPC examples (Matrix Multiplication, DGEMM, and Stream). We verified that the code compiled successfully for C++ and FORTRAN. We verified that the codes executed without any runtime errors, and that the code produced correct results for all codes.
Language
C++
Fortran
Python
Julia
C++
Fortran
Python
Julia
C++
Fortran
Python
Julia
Example Matrix Multiplication DGEMM Stream
Compilation x x x
Execution x x x x x
Correctness x ?
DifficultGoodEasyPoorDifficultGoodPythonC++FortranJulia
(a) Conjugate gradient
DifficultGoodEasyPoorDifficultGoodPythonC++FortranJulia
(b) Heat equation
DifficultGoodEasyPoorDifficultGoodPythonC++JuliaFortran
(c) Matrix Multiplication
DifficultGoodEasyPoorDifficultGoodPythonC++JuliaFortran
(d) DGEMM
DifficultGoodEasyPoorDifficultGoodPythonC++JuliaFortran
(e) STREAM Triad
Figure 1: Two-dimensional classification using the estimated schedule effort of the COCOMO model (easy vs difficult) and the quality of the code using compilation, execution, and correctness (poor vs good).
4.6 Programming efforts
Similar to the quality, the effort to write the code generated by a programmer is investigated. One way to estimate that effort is by using the Constructive Cost Model (COCOMO) [2, 19]. The COCOMO model does not take parallelism into account and ignores features like synchronization. Starting in the 90s, the HPC community discussed the development of a similar model for parallel and distributed applications. No such model has been proposed as of the time of this writing. One attempt was made to add parallel programming to the COCOMO II model [18].
We used the tool scc [^2] https://github.com/boyter/scc to estimate the human effort to produce the code. We use Tables 3,4 to quantify the quality of the software from poor to good. Figure 1 shows the estimated programming efforts and quality of the code. For the conjugate gradient the C++ code needed more effort. For the heat equation the effort was more balanced. For the matrix multiplication the C++ code required the most effort. For DGEMM and STREAM Triad the effort was more balanced.
5 Performance evaluation
For the performance evaluation, the following versions were used: g++/gfortran 13/14, Julia 1.11.3 and Python 3.11. Recall that we just fixed compilation and runtime errors in the code, but did not improve the parallel implementations. Figure 2(a) shows the scaling for the generated codes on AMD EPYC 7763 (x86) from one thread to 64 cores. We used $10,000,000$ nodes and changed the length of the domain $L$ to $100,000$. The C++ code and Python code scaled with the number of cores. The Fortran and Julia code showed some speedup from a single core to five cores. After that, the code did not benefit from additional cores. Figure 2(b) shows the scaling from one core to 64 cores on Intel Xeon Platinum 8358 (x86). The matrix size was $000\times 10,$ 000 × 10 , 000 with $100,000,000$ elements. Here, only the Julia, code scaled. Python, C++  and Fortran showed strange behavior. Figure 2(c) shows the DGEMM benchmark for matrices with 512, 1024, and 2048 rows and columns and a block width of 64 on Arm A64FX. These numbers were generated by the LLM. The Fortran code was very slow and the GFLOP/s were in the range of 0.02. The Python code reported around the same values for all sizes. The C++ and Julia code had increasing numbers. However, the Julia codes reported GFLOP/s for a BLAS implementation and these values were hundreds times bigger. Thus, all of the codes had a very poor performance. Figure 2(d) shows the performance for stream triad for $10^{4}$, $10^{5}$, and $10^{6}$ elements on ARM Grace Grace using 72 cores. Here, all code increased performance with the array size. For the Fortran code, we used opencoarrays with openmpi 4.1 as the coarray implementation.
$2^{0}$$2^{1}$$2^{2}$$2^{3}$$2^{4}$$2^{5}$$2^{6}$$2^{5}$$2^{8}$$2^{11}$$2^{14}$$2^{17}$# coresprocessed points per secondParallel heat equation (x86-AMD)
$2^{0}$$2^{1}$$2^{2}$$2^{3}$$2^{4}$$2^{5}$$2^{6}$$2^{14}$$2^{16}$$2^{18}$$2^{20}$$2^{22}$$2^{24}$# coresprocessed matrix entries per secondParallel Matrix Multiplication (x86-Intel)
51210242048002222444466660.10.10.10.10.20.20.20.20.40.40.40.40000000.30.30.30.30.30.30.30.30.30.30.30.32.52.52.52.54.54.54.54.54.14.14.14.1row and colnum sizeGFLOP/sDGEMM (A64FX)C++FortranPythonJulia
$00010001{,}$,50015001{,}5001 , 50024242424473473473473348348348348404040403003003003001,20012001{,}2001 , 2000.20.20.20.22.22.22.22.220202020525252527070707069696969# elementsGB/sStream triad (Grace Grace)C++FortranPythonJulia
Figure 2: Performance measurements: (2(a)) parallel heat equation solver on x86, (2(b)) parallel matrix multiplication on RISC-V, (2(c)) DGEMM on Arm A64FX, and (2(d)) stream triad on Arm Grace Hopper.
6 Discussion and Conclusion
The results of this study reveal both the potential and limitations the LLM model DeepSeek in HPC. By evaluating the performance of generated code across different benchmarks, we identified key insights regarding code quality, scalability, and parallel execution.
One important observation is that LLMs struggle to write scalable codes. In first two examples, both the C++ and Python codes showed consistent scaling with the increase in the number of cores, while the Fortran and Julia implementations showed limited scalability. This is consistent with prior research on HPC applications, which highlights the challenges of parallelism in certain programming languages. In example three, both C++ and Fortran scaled, but the Julia code could not be executed on RISC-V CPUs due to Julia not supporting this device. The Stream Triad clearly improved in performance with increasing array sizes for the all generated codes. Using COCOMO analysis, the C++  code required more effort in the conjugate gradient and matrix multiplication examples. While the effort was more balanced between all the languages for the heat equation solver and STREAM Triad tasks.
DeepSeek shows promise in generating code across several HPC benchmarks, however the performance of the generated code needs more improvement for HPC tasks. The results also highlight the difficulties in optimizing for parallel execution and memory efficiency when generating code with LLMs. Future work should focus on improving the parallelization and performance of the generated code for HPC tasks like DGEMM and Stream Triad. To conclude, when used alongside traditional programming methods, LLMs can significantly reduce the effort required for code generation, but there is still work to be done to reach the performance levels expected for high-performance applications.
In a future work we will study the performance for distributed applications, like MPI, acceleration cards, and abstraction layers (like Kokkos or SYCL).
Supplementary materials
The generated source code is available on GitHub [^3] https://github.com/NoujoudNader/AiCode_DeepSeek or Zenodo [^4] https://doi.org/10.5281/zenodo.14968599, respectively.
Acknowledgments
This work was supported by the U.S. Department of Energy through the Los Alamos National Laboratory. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 12345) LA-UR-25-22174

### v2

@@
 Their DeepSeek-V3 [20, 17] was trained in two months at a cost of \$5.6 million, significantly cheaper than similar models.
-On January 20th, they launched DeepSeek-R1 [13], a reasoning model exhibiting performance comparable to OpenAI’s O1 model, available for open research [9].
+In contrast, the development cost of ChatGPT-4 ranged between \$41 million and \$78 million [3].
+On January 20th 2025, they launched DeepSeek-R1 [16], a reasoning model exhibiting performance comparable to OpenAI’s O1 model, available for open research [12].
 DeepSeek-R1 is trained through large-scale reinforcement learning (RL) without supervised fine-tuning, demonstrating powerful reasoning behaviors [16].
@@
 The PLs are C++ for performance and simulation, Fortran for numerical and array optimizations, Python for flexibility and integration with other HPC languages, and Julia for scientific computing performance with a Python-like syntax.
-We present five examples to assess the model performance.
+We present five examples to assess the performance of the model.
 The first example in this study features a conjugate gradient solver that incorporates matrix and vector operations.
 The second example is a parallel 1D stencil-based heat equation solver [10], chosen for its relevance in testing parallel computational workflows and its widespread use in scientific computing.
-The third example is parallel matrix multiplication, a fundamental operation in scientific computing and machine learning, often accelerated using HPC systems to enhance computational speed.
+The third example is parallel matrix multiplication, a fundamental operation in scientific computing and machine learning, often accelerated by using HPC systems to enhance computational speed.
 The fourth example is parallel double-precision general matrix multiply (DGEMM), a widely used operation in numerical linear algebra that is essential for many scientific and engineering applications, testing the system’s computational efficiency for large-scale matrix operations.
@@
 LLMs have been widely used across various domains, including natural language processing (NLP) and visualization.
+Multiple efforts have also been made to benchmark LLMs for tasks such as code generation [9, 8, 6].
 However, their application in analyzing and optimizing HPC tasks remains challenging.
 In Table 1, we present the related works that use LLMs for HPC tasks.
-LLM4HPC [4] represents the first effort to adapt LLMs specifically for the HPC domain.
+LLM4HPC [5] represents the first effort to adapt an LLM specifically to the HPC domain.
 The LLM4HPC framework is specifically designed for HPC, showing success in code similarity analysis, parallelism detection, and OpenMP question-answering tasks.
-LLM4VV [20] is a fine-tuned model that leverages the capabilities of GPT-4 and CodeLlama (based on Llama-2) to successfully generate OpenACC directives.
+LLM4VV [23] is a fine-tuned model that uses the capabilities of GPT-4 and CodeLlama (based on Llama-2) to successfully generate OpenACC directives.
 HPC-GPT introduced by Ding et al.
@@
 Table 2: Queries for HPC tasks code generation, whether the language was C++, Fortran, Julia, Python
-Ex. Problem Query
+Ex. Problem Prompt
 1 Conjugate Gradient Solver Write a language code to solve the linear equation system using the conjugate gradient solver and validate it.
@@
 the Euler method.
-Example 3: In this example, we evaluate DeepSeek’s ability to generate parallel code for matrix multiplication, a fundamental operation in numerical simulations and machine learning.
-Matrix multiplication is a critical task in scientific computing, particularly in HPC, where large matrices are multiplied to generate results in parallel.
-The task involves writing efficient parallel code to split matrix multiplication across multiple threads or processors, optimizing for memory usage and computational speed.
+Example 3: In this example, we evaluate DeepSeek’s ability to generate parallel code for matrix multiplication, a general benchmarking task used to evaluate the basic performance and scalability of parallel code.
+It is also a fundamental operation in numerical simulations and machine learning.
+The task involves writing efficient parallel code that runs across multiple threads or processors, optimizing for memory usage and computational speed.
 The generated code is supposed to compute the following:
 $C(i,j)=C(i,j)+A(i,k)*B(k,j)\text{,}\quad A,B,C\in\mathbb{R}^{n\times n}\text{.}$ (5)
-Example 4: The fourth example is more general and focuses on DGEMM, a well known operation in scientific computing and HPC.
+Example 4: The fourth example is more general and focuses on DGEMM, which targets a more specialized application within the HPC domain.
 The objective here is to analyze the ability of DeepSeek to generate parallel code for performing DGEMM on large-scale matrices, with a particular focus on optimizing for performance using parallel computing techniques.
@@
 $C=\alpha\cdot A\cdot B+\beta\cdot C,\quad A,B,C\in\mathbb{R}^{n\times n},\alpha,\beta\in\mathbf{R}\text{.}$ (6)
+While Example 3 is useful for assessing general parallel performance, DGEMM provides a deeper evaluation of optimized performance for large-scale, high-precision operations.
 Example 5: The final example evaluates DeepSeek’s ability to implement the STREAM triad operation, which is commonly used in HPC benchmarks to measure memory bandwidth and data throughput.
@@
 Correctness x ?
-DifficultGoodEasyPoorDifficultGoodPythonC++FortranJulia
+DifficultGoodEasyPoorDifficultGoodpyC++FortranjlpyC++Fortranjl
 (a) Conjugate gradient
-DifficultGoodEasyPoorDifficultGoodPythonC++FortranJulia
+DifficultGoodEasyPoorDifficultGoodpyC++FortranjlpyC++Fortranjl
 (b) Heat equation
-DifficultGoodEasyPoorDifficultGoodPythonC++JuliaFortran
+DifficultGoodEasyPoorDifficultGoodpyC++jlFortran
 (c) Matrix Multiplication
-DifficultGoodEasyPoorDifficultGoodPythonC++JuliaFortran
+DifficultGoodEasyPoorDifficultGoodpyC++jlFortran
 (d) DGEMM
-DifficultGoodEasyPoorDifficultGoodPythonC++JuliaFortran
+DifficultGoodEasyPoorDifficultGoodpyC++jlFortran
 (e) STREAM Triad
 Figure 1: Two-dimensional classification using the estimated schedule effort of the COCOMO model (easy vs difficult) and the quality of the code using compilation, execution, and correctness (poor vs good).
+The blue values show the results from ChatGPT [9] and the black values the results for Deep Seek.
+The Python and Julia data points are tagged with the common file endings py and jl, respectively.
 4.6 Programming efforts
@@
 For the Fortran code, we used opencoarrays with openmpi 4.1 as the coarray implementation.
+Table 5 shows the peak performance for the parallel heat equation and matrix multiplication.
 $2^{0}$$2^{1}$$2^{2}$$2^{3}$$2^{4}$$2^{5}$$2^{6}$$2^{5}$$2^{8}$$2^{11}$$2^{14}$$2^{17}$# coresprocessed points per secondParallel heat equation (x86-AMD)
@@
 $00010001{,}$,50015001{,}5001 , 50024242424473473473473348348348348404040403003003003001,20012001{,}2001 , 2000.20.20.20.22.22.22.22.220202020525252527070707069696969# elementsGB/sStream triad (Grace Grace)C++FortranPythonJulia
-Figure 2: Performance measurements: (2(a)) parallel heat equation solver on x86, (2(b)) parallel matrix multiplication on RISC-V, (2(c)) DGEMM on Arm A64FX, and (2(d)) stream triad on Arm Grace Hopper.
+Figure 2: Performance measurements: (2(a)) parallel heat equation solver on x86-AMD, (2(b)) parallel matrix multiplication on x86-Intel, (2(c)) DGEMM on Arm A64FX, and (2(d)) stream triad on Arm Grace Hopper.
+Table 5: Peak performance for the parallel heat equation and parallel matrix multiplication as points processed per second and matrix entries processed per second, respectively.
+The number in parentheses are the number of cores at the peak performance.
+Example C++ Fortran Python Julia
+Heat equation 64 790 3386479033864\,790\,33864 790 338 (64) 604 750604750604\,750604 750 (15) 8 127 13781271378\,127\,1378 127 137 (64) 166 899166899166\,899166 899 (15)
+Matrix multiplication 192 661192661192\,661192 661 (60) 7 502 60775026077\,502\,6077 502 607 (25) 169 678169678169\,678169 678 (60) 1 225 91912259191\,225\,9191 225 919 (55)
 6 Discussion and Conclusion
@@
 This is consistent with prior research on HPC applications, which highlights the challenges of parallelism in certain programming languages.
-In example three, both C++ and Fortran scaled, but the Julia code could not be executed on RISC-V CPUs due to Julia not supporting this device.
+In example three, both Python and Fortran did not scale.
 The Stream Triad clearly improved in performance with increasing array sizes for the all generated codes.